OpenStudy (dehelloo):

-(-D) Wha does the minus sign outside the bracket do to the -D inside the bracket? 100% serious question...lol

Nnesha (nnesha):

when you multiply two same signs you get positive outcome.

Nnesha (nnesha):

+ times + = + - times - = +

Nnesha (nnesha):

and btw -(-D) is same as -1(-D) or you can rewrite it as \[\rm -*-D~~~ or ~~~~-1 *-D\]

rvc (rvc):

i have a tutorial on this check my profile

OpenStudy (mathmale):

Yes. Nnesha is right. What, then, is the (simplified) final result if you work on -(-D)? -(-D) = ?? (This question is for @dehelloo.)

OpenStudy (dehelloo):

D lol ?

OpenStudy (dehelloo):

Why a negative times a negative is a positive

OpenStudy (dehelloo):

I try to conceptualize it but still don;t understand why double neg equals positive.

OpenStudy (dehelloo):

-5 x 4 = -15 -5 + -5 + -5 +-5 or -5-5-5-5 -5 x -4 = 15 but (-5)- (-5) - (-5) - (-5) I just don;t understand why lol other than deducing that two negatives should equal a positive because -5(4+-4) = 0

OpenStudy (dehelloo):

Woops -20 lol

Directrix (directrix):

Think of -(-5) and being 0 - (-5). The rule of subtraction is to change the sign of the number being subtracted (called the subtrahend) and to add it to the first number. 0 - (-5) becomes 0 + (+5) = 0 + 5 = 5. Less mathematically but perhaps more interesting. Here's what a teacher once told me. "Negative numbers are living in a town with 'nice' numbers. Negative numbers are bad. A minus symbol appears when a person leaves town. Therefore, -(-D) means the bad person -D is leaving town. That is a good event. Ergo, -(-D) = +D."

OpenStudy (dehelloo):

How would you use the analogy for bad/good numbers for a singular substraction for -D-D? That there are bad people leaving bad people but it results in more bad people/numbers. Appreciate the answers though.

Directrix (directrix):

-D -D = - (D + D) by the distributive property - (D + D) = - (2D) - (2D) = 0 - (2D) Consider 2D as two good people 0 - (2D) = 0 +(-2D) = + (-2D) = -2D Two bad people came to town and that is a bad thing. Note that the explanation is bordering on absurdity.

OpenStudy (dehelloo):

Lol thanks for trying to help. Makes no sense but sort of does so I guess I'll just have to accept the law as fact.